# Optimizing Token Choice for Code Watermarking: A RL Approach

### v1

@@ 5.2 Robustness Analysis
Table 2: Robustness evaluation (%) under code modification attacks. Bold values indicate best performance.
@@
Figure 2: Performance for OpenCoder-8B-Instruct.
Figure 3: The training dynamics for CodeTracer-0.
As shown in Table 2, we evaluate CodeTracer’s robustness against DIPPER [10] paraphrasing and variable renaming attacks [13]. While all methods experience performance degradation under attacks, CodeTracer consistently outperforms competing approaches. Against DIPPER attacks, CodeTracer maintains the highest detection capability (AUROC 58.42%, TPR 14.31%). Similarly, under variable renaming attacks, CodeTracer demonstrates superior resilience (AUROC 73.36%, TPR 29.11%), indicating its effectiveness in real-world scenarios with code modifications.
@@ 5.3 Further Analysis
Scalability to Larger Models.
To verify the scalability or our approach, we evaluate CodeTracer on the larger OpenCoder-8B-Instruct model. Results are shown in Figure 3. CodeTracer achieves minimal code functionality degradation (71.77% Pass@1 vs. 72.04% for Base) while significantly outperforming other practical watermarking methods in detection capability (78.69% AUROC compared to 65.90% for WLLM and 54.21% for EXP-edit). These results demonstrate CodeTracer’s effectiveness and generalizability across different parameter scales.
Figure 4: Performance (%) comparison with variant
Dataset Performance Metrics
Method Pass@1 AUROC TPR
CodeTracer 60.82 82.95 46.34
CodeTracer-0 58.65 77.50 27.00
Training from Scratch.
To evaluate the effectiveness of pure reinforcement learning without prior knowledge, we introduce CodeTracer-0, a variant trained without SFT initialization. As shown in Figure 3, despite beginning with low detection rewards (approximately 0.3), the model consistently improves throughout RL training, eventually stabilizing around 0.8 for outcome-based watermark detection reward $R_{2}$, which is a proxy for z-score. Table 4 quantifies the final performance: CodeTracer-0 achieves reasonable results with Pass@1 of 58.65%, AUROC of 77.50%, and TPR of 27.00%, though not matching our fully-initialized CodeTracer. These results demonstrate that while good initialization enhances RL performance, our watermarking method remains effective even when trained entirely from scratch.
Figure 5: Computational Efficiency Comparison
Component Resource Metrics
Method Parameters Memory Inference
[…]
Baseline 1.5B Model 1,500M $\sim$3-6GB $\sim$500-800ms
Relative Overhead $10\% <$$15\% <$$0.02\%$
Computation Efficiency
[…]
@@ 6 Conclusion
In this paper, we presented CodeTracer, a novel adaptive code watermarking framework that addresses the challenges of watermarking LLM-generated code while preserving functionality. Our policy-driven approach integrates a watermark model with the base LLM to form a composite policy that dynamically decides when to watermark and which tokens to select. By leveraging reinforcement learning through GRPO, our approach intelligently navigates programming language constraints without requiring pre-existing watermarked examples. Our multi-component reward system balances statistical detectability with code integrity through execution feedback and token-level guidance. Evaluations demonstrate that CodeTracer significantly outperforms existing approaches in both detection accuracy and functionality preservation, offering a promising solution for identifying AI-generated code while maintaining its utility.
7 Limitations
While CodeTracer demonstrates significant advantages, certain limitations merit consideration. The reinforcement learning process for training the watermarking policy $\pi_{\theta}$, though optimized with rule-based rewards, can be computationally intensive, especially when applied to extensive codebases or a wide array of programming languages. The evolving nature of adversarial attacks also means that while CodeTracer enhances robustness, ongoing research will be crucial to maintain resilience against future sophisticated attempts at watermark evasion.
The dynamic, per-token decision-making by the watermarked policy $\pi_{\theta\oplus\phi}$, including its determination of green/red list partitions, could also introduce a slight inference latency during code generation compared to simpler, non-adaptive watermarking methods; the impact of this would depend on the specific complexity of the learned policy. Finally, while our use of Gumbel Soft Top-k effectively enables differentiable learning for discrete vocabulary partitioning, the characteristics of this approximation and its potential interaction with highly diverse or unusual coding constructs represent an area for continued detailed analysis.
@@ A.1 Watermark Generation
4:  Compute context window $\mathbf{ctx}=\text{concat}(\mathbf{x},\mathbf{\tilde{s}})[-(c):]$
@@
6:  Compute watermark decision $w=\mathbbm{1}_{w_{\phi}>0.5}$
@@
14:  Append $\tilde{s}_{t}$ to $\mathbf{\tilde{s}}$
15:end for
16:return  $\mathbf{\tilde{s}}$
@@ A.2 Watermark Detection
Algorithm 2 details our statistical watermark detection procedure. Following the framework of [9], we employ hypothesis testing with a key modification, that testing is performed selectively only at positions likely to contain watermarks.
Given a code sequence $\mathbf{s}$, the detector first reconstructs the vocabulary partitions and switch probabilities using the same watermark model from generation. The detection employs a one-sided hypothesis test using the z-statistic to determine whether green tokens appear more frequently than random chance would predict, with a positive z-score above the threshold indicating watermark presence.
Algorithm 2 CodeTracer Watermark Detection
@@
4:  Compute context window $\mathbf{ctx}=\mathbf{s}[\max(0,t-c):t]$
@@
12:   end if
13:   Add $t$ to positions_checked
14:  end if
15:end for
16: Compute z-score: $z=\frac{\text{green\_tokens}-\text{watermarked\_positions}\cdot\gamma}{\sqrt{\text{watermarked\_positions}\cdot\gamma\cdot(1-\gamma)}}$
17:if $z>\tau$ then
18:  return  “watermarked”
19:else
20:  return  “not watermarked”
21:end if
@@ B.1 SFT Initialization
While our reinforcement learning (RL) algorithm can train from random initialization, we find that supervised fine-tuning (SFT) provides a significantly more effective starting point. Since watermarked code is unavailable for direct supervision, we instead use entropy and next token prediction of clean code as a proxy. During SFT, the watermark model $\pi_{\phi}$ is trained on two complementary objectives: predicting watermark decisions $w_{\phi}$ to learn entropy distribution patterns across code structures, and optimizing logits $l_{\phi}$ to capture the underlying distribution of tokens in code sequences. This dual-objective approach establishes a strong foundation that encodes both code syntax patterns and appropriate watermarking signals before RL optimization begins.
@@ B.4 Reward Design for Trivial Solutions
where $\alpha>1$ (typically 2-5) creates a stronger penalty for red token selection when watermarking is active. This asymmetric design incentivizes the model to be more selective about when to apply watermarking ($w=1$), only doing so when it has high confidence in successfully biasing toward green tokens. For less certain positions, the model learns to disable watermarking ($w=0$), which receives a neutral reward. This approach significantly reduces false positive detection rates while maintaining high true positive rates.
@@ Appendix C Implementation Details
Table 3 presents the comprehensive hyperparameter configuration for our CodeTracer framework, categorized by functional components to facilitate reproducibility and highlight critical design decisions across our experimental setup. For paper writing, we use AI tools to refine the overall expression.
@@ C.1 Datasets
We evaluate on standard code generation benchmarks that span multiple programming languages and tasks. The primary evaluation uses HumanEval [3] and MBPP [2], which provide comprehensive Python programming challenges with associated test cases and reference implementations.
@@ C.2 Baselines
We compare against two categories of methods: post-hoc detection and active watermarking. Post-hoc detection methods preserve original generation and include zero-shot approaches: logp(x), LogRank [5], DetectGPT [18], and GPTZero [21]. Active watermarking methods include WLLM [9] and EXP-edit [11], which operate under the practical constraints. We also include SWEET [13] as a reference, though comparisons are not fair since it requires access to the original LLM and prompts for entropy calculation. We thoroughly tune the hyperparameters for each baseline following their setup.
@@
For the DetectGPT implementation, we used T5-3B as our model. Following the original DetectGPT paper and SWEET [18, 13] , we set the span length to 2 words and applied masking to 20% of the text. For each test, we generated 100 perturbations to ensure robust detection.
@@
For the SWEET baseline implementation, we conducted a systematic hyperparameter search for the entropy threshold, exploring values from 0.3 to 1.2 with increments of 0.3. Our experiments revealed optimal performance with an entropy threshold of 1.2 for the HumanEval dataset and 0.3 for MBPP. In detection staga, SWEET needs to use original LLM, prompt and complete generation sequence to compute the entropy. This requirement severely restricts their practical applicability in real-world scenarios. For practical deployment, an ideal solution would depend solely on the code snippet itself, as access to the generation context is typically unavailable. Thus, we only use SWEET as a reference model in our performance comparision.
@@
We evaluate EXP-edit [11] following the methodology of [13]. In our experiments, we set temperature=$0.2$ and top-p=$0.95$. We systematically explore block sizes ($20$ tokens), key sequence lengths ($100$), resample sizes ($50$ runs), and edit distance thresholds ($\gamma=0.0$). Through extensive parameter tuning, we determine the optimal configuration: key length $100$, block size $20$, $50$ sampling runs, and detection threshold $0.1$, which balances watermark detection reliability with code functionality.
@@ C.4 Training Details
We implement reinforcement learning training using GRPO with several customizations for the code watermarking task. Our training pipeline processes the mix of open-r1/verifiable-coding-problems-python dataset, HumanEval dataset and MBPP dataset. The motivation is to make sure model be familiar the distribution of these dataset instructions, which are representative code examples for code generation tasks.
The watermark policy model optimization uses a learning rate of $1.0\times 10^{-5}$ with a cosine scheduler including a minimum learning rate rate of 0.1. We implement a short warmup period (3% of total steps) to stabilize early training. For regularization, we set $\beta=0.0$ for the KL divergence term, allowing the watermark policy to explore freely while process-based rewards provide necessary constraints. The training process runs for 500 steps with a batch size of 8 per device and gradient accumulation every 4 steps. For each prompt, we generate 8 different code completions to provide robust training signals. The maximum prompt length is limited to 256 tokens, while completions can extend up to 1024 tokens to accommodate various code generation tasks. During inference, we use a temperature of 1.0 to maintain natural generation characteristics.

### v2

@@ 1.5B LLM $\pi_{\theta}$ 1,500M  3-6GB  500-800ms
Table 3: Robustness evaluation (%) under code modification attacks. Bold values indicate best performance. Settings follow Lee et al. (2023).
@@
Table 4: Ablation study on different reward components for CodeTracer training.
Method Pass@1 (%) AUROC (%) TPR (%)
CodeTracer 60.82 82.95 46.34
w/o $A_{2}$ 61.15 (+0.33) 75.11 (-7.84) 30.29 (-16.05)
w/o $A_{1}$ 60.34 (-0.48) 79.52 (-3.43) 34.91 (-11.43)
Table 5: Performance of model variants.
Method Pass@1 (%) AUROC (%) TPR (%)
CodeTracer-1 62.65 77.71 32.32
CodeTracer-2 60.82 82.95 46.34
Figure 3: Cross-language evaluation on Java and C++. CodeTracer achieves consistent cross-language performance.
Figure 4: Training dynamics for CodeTracer-1 with RL training.
Transferability to Larger Models. To verify the scalability of our approach, we train the watermark model on a smaller 1.5B LLM and directly apply it as a plug-in module to the larger OpenCoder-8B-Instruct without retraining. Results are shown in Figure 2. CodeTracer achieves minimal code functionality degradation (71.77% Pass@1 vs. 72.04% for Base) while significantly outperforming other watermarking methods in detection capability (78.69% AUROC compared to 65.90% for WLLM and 54.21% for EXP-edit). This demonstrates that our watermark model can serve as an effective plug-in module for larger models without requiring model-specific retraining.
Robustness Analysis. Table 5.1 evaluates CodeTracer’s robustness against DIPPER (Krishna et al., 2024) paraphrasing and variable renaming attacks (Lee et al., 2023). While all methods degrade under attacks, CodeTracer consistently outperforms baselines. Against DIPPER attacks, CodeTracer maintains superior detection (AUROC 58.42%, TPR 14.31%), and under renaming attacks demonstrates resilience (AUROC 73.36%, TPR 29.11%), indicating effectiveness in real-world scenarios.
Cross-Language Capabilities. Figure 4 shows CodeTracer’s cross-language performance on Java and C++. CodeTracer achieves strong detection while maintaining functionality comparable to baselines across languages, showing consistent effectiveness across different programming languages.
@@ 5.2 Further Analysis
Reward Components. Table 5.1 evaluates the impact of different reward components on CodeTracer’s performance. CodeTracer achieves optimal performance with both process and outcome rewards. Ablating either component leads to notable degradation, with removing the process reward causing larger detection drops while removing the outcome reward affects both performance.
Performance Tradeoff. Table 5.1 demonstrates the controllable trade-off between code functionality and watermark detectability. CodeTracer-1 uses solely RL training without supervised fine-tuning initialization, achieving higher functionality (Pass@1 62.65%) but lower detection performance (AUROC 77.71%). Figure 4 shows the training dynamics of CodeTracer-1, demonstrating consistent improvement throughout RL training. In contrast, CodeTracer-2 incorporates both supervised fine-tuning and RL, trading some functionality for significantly improved detection capability.
@@ 6 Conclusion
In this paper, we presented CodeTracer, a novel adaptive code watermarking framework that addresses the challenges of watermarking LLM-generated code while preserving functionality. Our policy-driven approach integrates a watermark model with the base LLM to form a composite policy that dynamically decides when to watermark and which tokens to select. By leveraging reinforcement learning through GRPO, our approach intelligently navigates programming language constraints without requiring pre-existing watermarked examples. Our multi-component reward system balances statistical detectability with code integrity through execution feedback and token-level guidance. Evaluations demonstrate that CodeTracer significantly outperforms existing approaches, offering a promising solution for identifying AI-generated code while maintaining the code functionality.
@@ A.1 Watermark Generation
4:  Compute context window $\mathbf{ctx}=\text{concat}(\mathbf{x},\mathbf{\tilde{s}})[-(c)\mathrel{\mathop{\ordinarycolon}}$]
@@
6:  Compute watermark decision $w=\mathbbm{1}_{w_{\phi}>0}$
@@
13:  Append $\tilde{s}_{t}$ to $\mathbf{\tilde{s}}$
14:return $\mathbf{\tilde{s}}$
Algorithm 1 CodeTracer Watermark Generation
@@ A.2 Watermark Detection
Algorithm 2 details our statistical watermark detection procedure. Following the framework of Kirchenbauer et al. (2023a), we employ hypothesis testing with a key modification, where testing is performed selectively only at positions likely to contain watermarks.
Given a code sequence $\mathbf{s}$, the detector first reconstructs the vocabulary partitions and switch probabilities using the same watermark model from generation. Detection employs a one-sided hypothesis test using the z-statistic to determine whether green tokens appear more frequently than random chance would predict, with a positive z-score above the threshold indicating watermark presence.
@@
14:  Compute context window
$\mathbf{ctx}=\mathbf{s}[\max(0,t-c)\mathrel{\mathop{\ordinarycolon}}t]$
@@
12:   Add $t$ to positions_checked
213: Compute z-score:
$z=\frac{\text{green\_tokens}-\text{watermarked\_positions}\cdot\gamma}{\sqrt{\text{watermarked\_positions}\cdot\gamma\cdot(1-\gamma)}}$
14:if $z>\tau$ then
15:  return “watermarked”
16:else
17:  return “not watermarked”
Algorithm 2 CodeTracer Watermark Detection
@@ B.1 SFT Initialization
While our reinforcement learning (RL) algorithm can train from random initialization, we find that supervised fine-tuning (SFT) provides a significantly more effective starting point. Since watermarked code is unavailable for direct supervision, we instead use entropy and next-token prediction of clean code as a proxy. During SFT, the watermark model $\pi_{\phi}$ is trained on two complementary objectives: predicting watermark decisions $w_{\phi}$ to learn entropy distribution patterns across code structures, and optimizing logits $l_{\phi}$ to capture the underlying token distribution in code sequences. This dual-objective approach establishes a strong foundation that encodes both code syntax patterns and appropriate watermarking signals before RL optimization begins.
@@ B.4 Reward Design for Trivial Solutions
where $\alpha>1$ (typically 2-5) creates a stronger penalty for red token selection when watermarking is active. This asymmetric design incentivizes the model to be more selective about when to apply watermarking ($w=1$), only doing so when it has high confidence in successfully biasing toward green tokens. For less certain positions, the model learns to disable watermarking ($w=0$), which receives a neutral reward. This approach significantly reduces false-positive detection rates while maintaining high true-positive rates.
@@ Appendix C Implementation Details
Table 6 presents the comprehensive hyperparameter configuration for our CodeTracer framework, categorized by functional components to facilitate reproducibility and highlight critical design decisions across our experimental setup.
@@ C.1 Datasets
We evaluate on standard code generation benchmarks that span multiple programming languages and tasks. The primary evaluation uses HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), which provide comprehensive Python programming challenges with associated test cases and reference implementations.
@@ C.2 Baselines
We compare against two categories of methods: post-hoc detection and active watermarking. Post-hoc detection methods preserve original generation and include zero-shot approaches: logp(x), LogRank (Gehrmann et al., 2019), DetectGPT (Mitchell et al., 2023), and GPTZero (Tian et al., 2023). Active watermarking methods include WLLM (Kirchenbauer et al., 2023a) and EXP-edit (Kuditipudi et al., 2023), which operate under the practical constraints. We also include SWEET (Lee et al., 2023), though comparisons are not fair since it requires access to the original LLM and prompts for entropy calculation. We adapt SWEET to use a practical setting for fair comparison with other baselines. We thoroughly tune the hyperparameters for each baseline following their setup.
@@
For the DetectGPT implementation, we used T5-3B as our model. Following the original DetectGPT paper and SWEET (Mitchell et al., 2023; Lee et al., 2023) , we set the span length to 2 words and applied masking to 20% of the text. For each test, we generated 100 perturbations to ensure robust detection.
@@
For the SWEET baseline implementation, we conducted a systematic hyperparameter search for the entropy threshold, exploring values from 0.3 to 1.2 with increments of 0.3. Our experiments revealed optimal performance with an entropy threshold of 1.2 and 0.9 for the HumanEval dataset and 0.3 and 0.6 for MBPP. In detection stage, SWEET needs to use original LLM, prompt and complete generation sequence to compute the entropy. This requirement severely restricts their practical applicability in real-world scenarios. For practical deployment, an ideal solution would depend solely on the code snippet itself, as access to the generation context is typically unavailable. Thus, we adapt SWEET to use fixed context windows (vs. full generation history and original LLM) for practical and fair comparison with other baselines. We select two entropy thresholds for SWEET to demonstrate the trade-off between watermark detectability and code functionality.
@@
We evaluate EXP-edit (Kuditipudi et al., 2023) following the methodology of (Lee et al., 2023). In our experiments, we set temperature=$0.2$ and top-p=$0.95$. We systematically explore block sizes ($20$ tokens), key sequence lengths ($100$), resample sizes ($50$ runs), and edit distance thresholds ($\gamma=0.0$). Through extensive parameter tuning, we determine the optimal configuration: key length $100$, block size $20$, $50$ sampling runs, and detection threshold $0.1$, which balances watermark detection reliability with code functionality.
@@ C.4 Training Details
We implement reinforcement learning training using GRPO with several customizations for the code watermarking task. Our training pipeline processes the mix of prompts from open-r1/verifiable-coding-problems-python dataset, HumanEval dataset and MBPP dataset. The motivation is to make sure model be familiar the distribution of these dataset instructions, which are representative code examples for code generation tasks.
The watermark policy model optimization uses a learning rate of $1.0\times 10^{-5}$ with a cosine scheduler including a minimum learning rate of 0.1. We implement a short warmup period (3% of total steps) to stabilize early training. For regularization, we set $\beta=0.0$ for the KL divergence term, allowing the watermark policy to explore freely while process-based rewards provide necessary constraints. The training process runs for 500 steps with a batch size of 8 per device and gradient accumulation every 4 steps. For each prompt, we generate 8 different code completions to provide robust training signals. The maximum prompt length is limited to 256 tokens, while completions can extend up to 1024 tokens to accommodate various code generation tasks. During inference, we use a temperature of 1.0 to maintain natural generation characteristics.
@@ Appendix D Security Discussion
Appendix E Use of Large Language Models
Large Language Models (LLMs) were used solely as a general-purpose writing assistant to check grammar and improve expression clarity throughout this paper. The LLMs did not contribute to research ideation, methodology development, experimental design, or generation of novel content. All technical contributions, insights, and findings presented in this work are entirely the result of the authors’ original research.